One pipe takes 30 minutes to fill a tank. After it has been running for 10 minutes, it is shut off. A second pipe is then opened and it finishes the tank in 15 minutes. How long would it take the second pipe alont to fill the tank?
Answer by nerdybill(7384)

One pipe takes 30 minutes to fill a tank. After it has been running for 10 minutes, it is shut off. A second pipe is then opened and it finishes the tank in 15 minutes. How long would it take the second pipe alont to fill the tank?
.
Let x = time (mins) it takes second pipe to fill the tank alone
then
10(1/30) + 15(1/x) = 1
multiplying both sides by 30x:
10(x) + 15(30) = 30x
10x + 450 = 30x
450 = 20x
22.5 mins = x